# Think2SQL: Reinforce LLM Reasoning Capabilities for Text2SQL

### v1

@@ Abstract
This paper investigates to what extent LLM reasoning capabilities influence their Text2SQL performance on four benchmark datasets. To this end, it considers the following LLM settings: (1) ZSL, including general-purpose reasoning or not; (2) SFT, with and without task-specific reasoning traces; (3) RL, leveraging execution accuracy as primary reward function; (4) SFT+RL, i.e, a two-stage approach that combines SFT and RL.
The results show that general-purpose reasoning under ZSL proves to be ineffective in tackling complex Text2SQL cases. Small LLMs benefit from SFT with reasoning much more than larger ones, bridging the gap of their (weaker) model pretraining. RL is generally beneficial across all tested models and datasets, particularly when SQL queries involve multi-hop reasoning and multiple tables. Small LLMs with SFT+RL excel on most complex datasets thanks to a strategic balance between generality of the reasoning process and optimization of the execution accuracy. Thanks to RL, the 7B Qwen-Coder-2.5 model performs on par with 100+ Billion ones on the Bird dataset.
@@ 1 Introduction
Reinforcement Learning (RL). We tailor RL to the Text2SQL task. The LLM repetitively performs actions consisting of shortlisting the best SQL query to solve the input question among a predefined set of candidates. We adopt Group-Relative Policy Optimization (GRPO) [41] and reward the LLM answers by using the established execution accuracy metric [56].
@@
To answer RQ2, we compare the performance of LLMs under (1) ZSL with reasoning vs. (2) SFT with reasoning vs. (3) RL vs. (4) SFT+RL. The goal is to compare different strategies to incorporate reasoning capabilities in LLM training for Text2SQL, with particular attention to the model performance achieved on complex SQL patterns.
To answer RQ3, we analyze the LLMs’ performance across diverse datasets, ranging from general-purpose to domain-specific content. The purpose is to clarify whether reasoning is beneficial to achieve model generality across different datasets and domains while preserving the overall accuracy of the SQL query generator.
@@ 2.3 Text2SQL
The Text2SQL task consists in mapping natural language question on databases to executable SQL code. Let ${{\boldsymbol{x}}}\in\mathcal{X}$ denote a natural language input (e.g., a user question), and let ${{\boldsymbol{y}}}\in\mathcal{Y}_{\text{sql}}$ denote a corresponding structured output in SQL syntax. The output space $\mathcal{Y}_{\text{sql}}$ comprises syntactically valid SQL codes consistent with a given database schema $\mathcal{S}$, which specifies the collection of relational tables, attributes, and their types. In addition to the schema, auxiliary context $\mathcal{M}$ may be provided. This includes task-specific metadata such as database descriptions, natural language annotations, examples of prior queries, or column-level summaries.
@@
The model $\pi_{{{\boldsymbol{\theta}}}}$ defines a conditional distribution over the SQL code ${{\boldsymbol{y}}}=(y_{1},\ldots,y_{T})$ given the input ${{\boldsymbol{x}}}$ and the schema context:
@@
For each prompt ${{\boldsymbol{x}}}$, the target may be a set of logically equivalent parses $\mathcal{Y}^{*}({{\boldsymbol{x}}})\subseteq\mathcal{Y}_{\text{sql}}$, all yielding identical execution results. Learning may proceed by optimizing the marginal log-likelihood over this set or by selecting a canonical representative from $\mathcal{Y}^{*}({{\boldsymbol{x}}})$ during training [58, 50, 43]. In this study, the latter approach is adopted, and the model is trained to predict a single SQL code ${{\boldsymbol{y}}}^{*}\in\mathcal{Y}^{*}({{\boldsymbol{x}}})$ for each input ${{\boldsymbol{x}}}$.
@@ 3 Methodology
Supervised Fine-Tuning (SFT). In the SFT approach, the model was trained on the curated dataset described in Section 3.1. The training objective was to minimize the cross-entropy loss between the predicted SQL code and the ground truth SQL code. Reasoning traces were included as additional input to guide the model in understanding the logical steps required to generate the correct SQL.
Reinforcement Learning (RL). For RL, we used execution accuracy as the primary reward signal. The model was fine-tuned using the GRPO algorithm, where the reward was computed based on the correctness of the generated SQL code’s execution results. To encourage the generation of reasoning traces, we also included a secondary reward signal based on syntatical checks. Details on the rewards are provided in Section 3.2.
@@ 3.1 SFT Dataset Creation
The BIRD training set comprises 9,428 data points derived from 69 heterogeneous databases spanning 37 professional domains, including blockchain, healthcare, education, and hockey. Each data point consists of a natural language (NL) question, a corresponding SQL code, and supplementary evidence. The evidence serves as additional context to resolve ambiguities in the schema or NL questions.
@@
Synthetic Data Annotation. To enhance the dataset with reasoning traces, we utilized the DeepSeek-R1 model [10] along with its system prompt. The prompt used for synthetic annotation is detailed in Figure 1. The hyperparameters optimized for reasoning tasks were set to a temperature of 0.7 and a top-p of 0.95, following established best practices in the field [48, 32]. This configuration ensured high-quality annotations that explicitly outline reasoning steps, enriching the dataset and enabling the model to better comprehend the logical processes underlying SQL generation.
@@ 3.2 Rewards for Reinforcement Learning
In reinforcement learning, reward signals are crucial for guiding the model’s learning process [17, 53, 11]. Execution accuracy, the primary reward for Text2SQL, measures the correctness of generated SQL by comparing it to the ground truth. However, its binary nature poses challenges for RL optimization, especially for smaller LLMs, as rewards often remain zero unless the SQL is exactly correct. To encourage the model’s reasoning process, we introduce an auxiliary reward signal that evaluates the correctness of the reasoning trace based on the appropriate use of reasoning tags [10]. Additionally, to mitigate reward hacking, a penalty is applied when reasoning tokens are redundantly or excessively repeated within the reasoning trace.
@@
Execution Accuracy (EX). Execution accuracy [56, 26] evaluates the correctness of the generated SQL by comparing its execution results against those of the ground truth SQL on the target database. This metric provides a binary reward, assigning a full score only when the two execution results match exactly, row by row. While straightforward and reliable, execution accuracy has a notable limitation: it does not account for partially correct results, such as queries that produce subsets or supersets of the expected results, which can hinder the learning process in reinforcement learning scenarios.
@@
Format Reward. The format reward [10] incentivizes the model to adhere to a predefined output structure, such as the use of <think> and <answer> tags. Outputs that conform to this structure receive a reward boost, promoting clarity and consistency. This is a sparse reward that activates only when both the opening and closing tags for reasoning and answers are correctly positioned. The reward value is 1 if the tags are correctly formatted; otherwise, it is 0.
Tag Count Reward. To address reward hacking, where reasoning traces include unnecessary or excessive tags, we introduce a tag count reward. This reward penalizes the model for generating reasoning traces with redundant tags. The reward is 1 if the number of tags is exactly two (one opening and one closing tag); otherwise, it is 0.
@@
The final reward signal is computed as a weighted sum of the individual rewards. The weights were carefully chosen to ensure a balanced contribution from each reward component while maintaining a total score of 1. This design prevents training instability caused by excessively high rewards. The final reward is computed as follows:
$R=0.85\cdot R_{\text{EX}}+0.10\cdot R_{\text{Format}}+0.05\cdot R_{\text{Tag Count}},\qquad R\in[0,1]$ (7)
@@
Table 1: Performance comparison of open-source and proprietary models on the Bird Dev dataset for the Text2SQL task. All models were evaluated with a temperature setting of 0.7 and a top_p value of 0.95. Llama models correspond to version 3.1 and Turbo means the model is quantized 8bit. Think2SQL-3B and Think2SQL-7B denote the models trained exclusively with RL.
@@
Think2SQL-3B $0.569$ $0.381$ $0.196$ $0.485$
Think2SQL-7B $0.619$ $0.463$ $0.406$ 0.552
4 Reasoning for Text2SQL
4.1 Experiment Setup
  • RQ1: Does reasoning improve the performance of Text2SQL models?
  • RQ3: How do models generalize to unseen databases?
We evaluate our models on the BIRD development set, which consists of $1,530$ instances (Simple #$\#924$, Medium #$\#461$ and Challenging #$\#143$), as the test set is not publicly available. To assess model robustness and generalization, we also evaluate on the SPIDER dataset [56], a widely recognized benchmark for Text-to-SQL tasks, along with its challenging variants Spider-Syn [14] and Spider-DK [15]. Spider-Syn tests robustness to paraphrased questions by introducing schema-related synonyms, while Spider-DK evaluates the model’s ability to incorporate domain knowledge by modifying both natural language questions and corresponding SQL queries to include implicit relationships or background knowledge not explicitly stated in the schema. For all evaluations, we report Execution Accuracy as the primary metric.
Training setup. To answer the posed research questions, we trained multiple models using different training strategies. Our experiments are based on the Qwen-Coder-2.5 model family [21], focusing specifically on the 3B and 7B variants.
@@ 4.2 Main Results
The Table 1 presents a performance comparison of various open-source and proprietary models on the Bird Dev dataset for the Text2SQL task, categorized by model size and reasoning capability. Among open-source models, the Qwen2.5-Coder-32B achieved the highest weighted average accuracy (0.553) in the 10-100B category, while Llama-405B-Turbo led the $>$100B category with a weighted average of 0.560. Proprietary models also performed competitively, with gpt-4o-2024-08-06 achieving a weighted average of 0.541. The Think2SQL models demonstrate strong performance across all evaluation categories, particularly on challenging examples. Think2SQL-7B achieves a weighted average score of 0.552, ranking second overall among all models tested—outperforming both open-source and closed-source models of significantly larger size, including Llama-70B-Turbo and multiple GPT-4 variants. Notably, it achieves the highest score on the Challenging subset (0.406), indicating superior reasoning and generalization abilities. Think2SQL-3B also performs competitively, with a weighted average of 0.485, surpassing all models below 10B parameters and several larger models, such as QwQ-32B and DeepSeek-Llama-70B.
When directly compared to their non-reasoning counterparts of similar size, the Think2SQL models consistently outperform them. Think2SQL-3B exceeds Qwen2.5-Coder-3B in all difficulty categories, with a weighted average of 0.485 ($+0.10$). Likewise, Think2SQL-7B outperforms Qwen2.5-Coder-7B across the board, particularly on moderate and challenging instances-0.388 vs. 0.463 ($+0.07$) and 0.294 vs 0.406 ($+0.11$), respectively. These improvements highlight the effectiveness of incorporating reasoning in training and underscore the competitive edge of our models even when compared to state-of-the-art baselines with similar architectures and parameter counts.
@@
Table 2: Ablation over different training strategies
Model Simple Medium Challenging AVG
@@
Qwen2.5-Coder-3B-SFTNO-THINK $0.510$ $0.325$ $0.224$ $0.427$
@@
Qwen2.5-Coder-3B-SFT-RLEX $0.560$ $0.370$ $0.343$ $0.482$
@@
Qwen2.5-Coder-7B-SFT $0.548$ $0.403$ $0.287$ $0.479$
Qwen2.5-Coder-7B-RLEX 0.619 0.463 0.406 0.552

### v2

@@ Abstract
This paper investigates to what extent LLM reasoning capabilities influence their Text2SQL performance on four benchmark datasets. To this end, it considers the following LLM settings: (1) ZSL, including general-purpose reasoning or not; (2) SFT, with and without task-specific reasoning traces; (3) RL, exploring the use of different rewarding functions, both the established EXecution accuracy (EX) and a mix with fine-grained ones that also account the precision, recall, and cardinality of partially correct answers; (4) SFT+RL, i.e, a two-stage approach that combines SFT and RL.
The results show that general-purpose reasoning under ZSL proves to be ineffective in tackling complex Text2SQL cases. Small LLMs benefit from SFT with reasoning much more than larger ones, bridging the gap of their (weaker) model pretraining. RL is generally beneficial across all tested models and datasets, particularly when SQL queries involve multi-hop reasoning and multiple tables. The use of the fine-grained metrics turns out to be the most effective RL strategy.
Small LLMs with SFT+RL excel on most complex datasets thanks to a strategic balance between generality of the reasoning process and optimization of the execution accuracy. Thanks to RL and the novel text2SQL rewards, the 7B Qwen-Coder-2.5 model performs on par with 400+ Billion ones (including gpt-4o) on the Bird dataset.
@@ 1 Introduction
Reinforcement Learning (RL). We tailor RL to the Text2SQL task. The LLM repetitively performs actions consisting of shortlisting the best SQL query to solve the input question among a predefined set of candidates. We adopt Group-Relative Policy Optimization (GRPO) [45] and explore the use of different rewarding functions encompassing both the established EXxecution accuracy (EX) metric [61] and a mix of fine-grained instance-level metrics [39, 39] that also account the precision, recall, and cardinality of partially correct Text2SQL answers.
@@
To answer RQ2, we compare the performance of LLMs under (1) ZSL with reasoning vs. (2) SFT with reasoning vs. (3) RL trained with EX vs. (4) RL trained with the novel Text2SQL rewards vs. (5) SFT+RL. The goal is to compare different strategies to incorporate reasoning capabilities in LLM training for Text2SQL, with particular attention to the model performance achieved on complex SQL patterns.
To address RQ3, we investigate the performance of LLMs trained with different RL reward functions. Specifically, we compare RL based on the traditional EX metric—the standard objective for most Text2SQL models, with recently proposed fine-grained instance-based metrics from the QATCH testing benchmark [39, 40], namely Cell precision, Cell recall, and Tuple cardinality. The key motivation is to address the shortcomings of EX, which acts as a sparse reward signal [34], failing to provide meaningful feedback when the model partially captures correct logical forms or schema relationships but otherwise receives no reward.
Leveraging RL with the newly introduced fine-grained metrics, the 7B Qwen-Coder-2.5 model achieves a 8.5% improvement over its base model, while the 3B variant achieves an 11.8% gain. Notably, the 7B model obtains the best performance among the evaluated models, surpassing even models with over 400 billion parameters. The 7B Qwen-Coder-2.5 model is publicly available on Hugging Face [^1] https://huggingface.co/simone-papicchio/Think2SQL-7B.
To answer RQ4, we analyze the LLMs’ performance across diverse datasets, ranging from general-purpose to domain-specific content. The purpose is to clarify whether reasoning is beneficial to achieve model generality across different datasets and domains while preserving the overall accuracy of the SQL query generator.
@@ 2.3 Text2SQL
The Text2SQL task consists in mapping natural language question on databases to executable SQL query. Let ${{\boldsymbol{x}}}\in\mathcal{X}$ denote a natural language input (e.g., a user question), and let ${{\boldsymbol{y}}}\in\mathcal{Y}_{\text{sql}}$ denote a corresponding structured output in SQL syntax. The output space $\mathcal{Y}_{\text{sql}}$ comprises syntactically valid SQL queries consistent with a given database schema $\mathcal{S}$, which specifies the collection of relational tables, attributes, and their types. In addition to the schema, auxiliary context $\mathcal{M}$ may be provided. This includes task-specific metadata such as database descriptions, natural language annotations, examples of prior queries, or column-level summaries.
@@
The model $\pi_{{{\boldsymbol{\theta}}}}$ defines a conditional distribution over the SQL query ${{\boldsymbol{y}}}=(y_{1},\ldots,y_{T})$ given the input ${{\boldsymbol{x}}}$ and the schema context:
@@
For each prompt ${{\boldsymbol{x}}}$, the target may be a set of logically equivalent parses $\mathcal{Y}^{*}({{\boldsymbol{x}}})\subseteq\mathcal{Y}_{\text{sql}}$, all yielding identical execution results. Learning may proceed by optimizing the marginal log-likelihood over this set or by selecting a canonical representative from $\mathcal{Y}^{*}({{\boldsymbol{x}}})$ during training [63, 55, 47]. In this study, the latter approach is adopted, and the model is trained to predict a single SQL query ${{\boldsymbol{y}}}^{*}\in\mathcal{Y}^{*}({{\boldsymbol{x}}})$ for each input ${{\boldsymbol{x}}}$.
@@ 3 Methodology
Supervised Fine-Tuning (SFT). In the SFT approach, the model was trained on the curated dataset described in Section 3.1. The training objective was to minimize the cross-entropy loss between the predicted SQL query and the ground truth SQL query. Reasoning traces were included as additional input to guide the model in understanding the logical steps required to generate the correct SQL.
Reinforcement Learning (RL). For RL, we used execution accuracy as starting point and introduce new reward for Text2SQL. The model was fine-tuned using the GRPO algorithm, where the reward was computed based on the correctness of the generated SQL query’s execution results. To encourage the generation of reasoning traces, we also included a secondary reward signals based on syntatical checks. Details on the rewards are provided in Section 3.2.
@@ 3.1 SFT Dataset Creation
The BIRD training set comprises 9,428 data points derived from 69 heterogeneous databases spanning 37 professional domains, including blockchain, healthcare, education, and hockey. Each data point consists of a natural language (NL) question, a corresponding SQL query, and supplementary evidence. The evidence serves as additional context to resolve ambiguities in the schema or NL questions.
@@
Synthetic Data Annotation. To enhance the dataset with reasoning traces, we utilized the DeepSeek-R1 model [10] along with its system prompt. The prompt used for synthetic annotation is detailed in Figure 1. The hyperparameters optimized for reasoning tasks were set to a temperature of 0.7 and a top-p of 0.95, following established best practices in the field [52, 35].
@@ 3.2 Rewards for Reinforcement Learning
In reinforcement learning, reward signals are crucial for guiding the model’s learning process [19, 58, 11]. Execution accuracy, the primary reward for Text2SQL, measures the correctness of generated SQL by comparing it to the ground truth. However, its binary nature poses challenges for RL optimization, especially for smaller LLMs, as rewards often remain zero unless the SQL is exactly correct. To address this limitation, we integrate QATCH [39, 40], an advanced benchmarking framework designed for the automated evaluation of Text2SQL tasks. For the purposes of this study, we employed three primary QATCH metrics: Cell Precision, Cell Recall, and Tuple Cardinality.
To encourage the model’s reasoning process, we introduce the Format reward that evaluates the appropriate use of reasoning tags [10]. Additionally, to mitigate reward hacking, the Tag count reward penalize the reward when reasoning tokens are redundantly or excessively repeated within the reasoning trace.
Let $\mathcal{T}$ and $\mathcal{T}_{\text{pred}}$ denote the execution results of the target SQL query and the predicted SQL query, respectively, each represented as a set of tuples, where each tuple comprises a set of cell values. The reward signals utilized in this study are outlined below:
Execution Accuracy (EX). EX [61, 28] evaluates whether the execution of the target SQL query matches the execution of the predicted SQL query. It is defined as:
$R_{\text{EX}}=\begin{cases}1&\text{if }\mathcal{T}=\mathcal{T}_{\text{pred}}\\ 0&\text{otherwise}\end{cases},\quad R_{\text{EX}}\in\{0,1\}$ (7)
@@
This metric provides a binary reward, assigning a full score only when the two execution results match exactly, row by row. While straightforward and reliable, execution accuracy does not account for partially correct results, which can hinder the learning process in RL.
Cell Precision (CP). CP is the fraction of table cells in $\mathcal{T}_{\text{pred}}$ that are in the target $\mathcal{T}$. The higher the score, the more predicted cells are in the target.
$R_{\text{CP}}=\frac{|\{cells\mid cells\in\mathcal{T}\cap\mathcal{T}_{\text{pred}}\}|}{|\{cells\mid cells\in\mathcal{T}\}|},\quad R_{\text{CP}}\in[0,1]$ (8)
This metric allows for partial credit when the predicted SQL query execution contains some requested cells but also includes incorrect ones. Considering the target query SELECT Name FROM Player; and the predicted query SELECT Name, Surname FROM Player; in this case CP is $0.5$ because the predicted SQL query execution contains the correct cells from the column Name and incorrect ones from Surname. However, it does not consider whether all the requested cells in $\mathcal{T}$ are present in the SQL query - measured by Cell Recall. It is worth noticing that when EX is $1$ also CP is $1$.
Cell Recall (CR). CR is the fraction of table cells in $\mathcal{T}$ that are present in $\mathcal{T}_{\text{pred}}$. The higher the score, the more target cells are included in the prediction.
$R_{\text{CR}}=\frac{|\{cells\mid cells\in\mathcal{T}\cap\mathcal{T}_{\text{pred}}\}|}{|\{cells\mid cells\in\mathcal{T}_{\text{pred}}\}|},\quad R_{\text{CR}}\in[0,1]$ (9)
This metric allows a partial reward in case the predicted SQL query does not contain all the requested cell in the target query. Considering the target query SELECT Name, Surname FROM Player; and the predicted query SELECT Name, Surname FROM Player; then CP is $1$ but CR is $0.5$ because the predicted SQL query execution contains the correct cells from the column Name but not from Surname. It is worth noticing that when EX is $1$ CR is $1$ as well.
Tuple Cardinality (TC). TC is defined as the ratio between the number of tuples in $\mathcal{T}_{\text{pred}}$ and the number of tuples in $\mathcal{T}$. The $\min$ function is used to ensure $\text{TC}\in[0,1]$. TC captures output cardinality only, ignoring schema and cell values. Thus, it should be considered alongside CP and CR for a fuller view of model performance. The TC reward is defined as:
$R_{\text{TC}}=\min\left(\frac{|\mathcal{T}|}{|\mathcal{T}_{\text{pred}}|},\frac{|\mathcal{T}_{\text{pred}}|}{|\mathcal{T}|}\right),\quad R_{\text{TC}}\in[0,1]$ (10)
This metric is necessary because CP and CR are computed based on the intersection of cell values, which may overlook differences in output size when the number of cells is not critical. For example, consider the target query SELECT DISTINCT Name FROM Player; and the predicted query SELECT DISTINCT Name FROM Player;. In this case, both CP and CR equal $1$ due to identical cell values, yet the cardinality of the cells is different. Thus, the TC metric is essential for capturing this discrepancy.
Format Reward (FR). The FR [10] incentivizes the model to adhere to a predefined output structure, such as the use of <think> and <answer> tags.
$R_{\text{FR}}=\begin{cases}1&\text{if }\pi_{{{\boldsymbol{\theta}}}}({{\boldsymbol{x}}})\text{ matches }\texttt{{\small<think/>.*?</think>s*<answer>.*?</answer>}}\\ 0&\text{otherwise}\end{cases},\quad R_{\text{FR}}\in\{0,1\}$ (11)
@@
This is a sparse reward that activates only when both the opening and closing tags for reasoning and answers are correctly positioned. The reward value is 1 if the tags are correctly formatted; otherwise, it is 0.
[…]
$R_{\text{TCR}}=0.25\cdot\sum_{t}\mathds{1}(Count(\pi_{{{\boldsymbol{\theta}}}}({{\boldsymbol{x}}}),t)=1),\quad R_{\text{TCR}}\in\{0,0.25,0.50,0.75,1.0\}$ (12)
@@
Final Reward. The final reward signal is computed as a weighted sum of the individual rewards. The weights were carefully chosen to ensure a balanced contribution from each reward component while maintaining a total score of 1. This design prevents training instability caused by excessively high rewards. In addition, since CP, CR, and TC must be seen together to provide a complete picture of the model’s performance, we decided to use the average of these three metrics $R_{\text{QATCH}}=\mathbb{E}[R_{\text{CP}},R_{\text{CR}},R_{\text{TC}}]$.
Let $R_{\text{text2SQL}}$ be $R_{\text{EX}}$ or $R_{\text{QATCH}}$, with the combination left for future study. The final reward is computed as follows:
$R=0.85\cdot R_{\text{text2SQL}}+0.10\cdot R_{\text{Format}}+0.05\cdot R_{\text{Tag Count}},\qquad R\in[0,1]$ (13)
This weights are selected to ensure that the execution accuracy and QATCH metrics are the primary focus of the training process, while still encouraging the model to produce well-structured outputs with appropriate reasoning traces.
4 Reasoning for Text2SQL
4.1 Experiment Setup
  • RQ1: Does reasoning improve the performance of Text2SQL models?
  • RQ3: Is EX the most effective reward function for Text2SQL RL?
  • RQ4: How do models generalize to unseen databases?
We evaluate our models on the BIRD development set, which consists of $1,530$ instances (Simple #$\#924$, Medium #$\#461$ and Challenging #$\#143$), as the test set is not publicly available. To assess model robustness and generalization, we also evaluate on the SPIDER dataset [61], a widely recognized benchmark for Text-to-SQL tasks, along with its challenging variants Spider-Syn [16] and Spider-DK [17]. Spider-Syn tests robustness to paraphrased questions by introducing schema-related synonyms, while Spider-DK evaluates the model’s ability to incorporate domain knowledge by modifying both natural language questions and corresponding SQL queries to include implicit relationships or background knowledge not explicitly stated in the schema. For all evaluations, we report Execution Accuracy as the primary metric and use LightEval [15] as base evaluation framework.
Training setup. To answer the posed research questions, we trained multiple models using different training strategies using as training framework Open-R1 [12, 53]. Our experiments are based on the Qwen-Coder-2.5 model family [23], focusing specifically on the 3B and 7B variants.
@@ 4.1 Experiment Setup
Table 1: Performance comparison of open-source and proprietary models on the Bird Dev dataset for the Text2SQL task. All models were evaluated with a temperature setting of 0.7 and a top_p value of 0.95. Llama models correspond to version 3.1 and Turbo means the model is quantized 8bit. Think2SQL-3B and Think2SQL-7B denote the Qwen2.5-Coder models trained exclusively with RL and $R_{\text{QATCH}}$.
@@
Think2SQL-3B $0.574$ $0.403$ $0.336$ $0.500$
Think2SQL-7B $0.628$ $0.482$ $0.385$ 0.561
@@ 4.2 Main Results
The Table 1 presents a performance comparison of various open-source and proprietary models on the Bird Dev dataset for the Text2SQL task, categorized by model size and reasoning capability. Among open-source models, the Qwen2.5-Coder-32B achieved the highest weighted average accuracy (0.553) in the 10-100B category, while Llama-405B-Turbo led the $>$100B category with a weighted average of 0.560. Proprietary models also performed competitively, with gpt-4o-2024-08-06 achieving a weighted average of 0.541. The Think2SQL models demonstrate strong performance across all evaluation categories, particularly on challenging examples. Think2SQL-7B achieves a weighted average score of 0.561, ranking first among all models tested—outperforming both open-source and closed-source models of significantly larger size, Notably, it achieves the highest score on the Challenging subset (0.385), indicating superior reasoning and generalization abilities. Think2SQL-3B also performs competitively, with a weighted average of 0.500, surpassing all models below 10B parameters and several larger models, such as QwQ-32B and DeepSeek-Llama-70B.
When directly compared to their non-reasoning counterparts of similar size, the Think2SQL models consistently outperform them. Think2SQL-3B exceeds Qwen2.5-Coder-3B in all difficulty categories, with a weighted average of 0.500 ($+0.12$). Likewise, Think2SQL-7B outperforms Qwen2.5-Coder-7B across the board, particularly on moderate and challenging instances-0.388 vs. 0.482 ($+0.09$) and 0.294 vs 0.385 ($+0.09$), respectively. These improvements highlight the effectiveness of incorporating reasoning in training and underscore the competitive edge of our models even when compared to state-of-the-art baselines with similar architectures and parameter counts.
@@
Table 2: Ablation over different training strategies. In bold are the best results for each model size. The RLEX and RLQATCH are the RL training strategies with EX and QATCH, respectively. The SFTNT is the SFT training strategy without reasoning tokens. In parenthesis, we show the relative improvement over the base model. The score is EX, higher is better.
Model Simple Medium Challenging Weighted AVG
@@
Qwen2.5-Coder-3B-SFTNT $0.510$ $0.325$ $0.224$ $0.427$
@@
Qwen2.5-Coder-3B-RLQATCH 0.574 (+10.5%) 0.403 (+13.6%) $0.336$ 0.500 (+11.8%)
Qwen2.5-Coder-3B-SFT-RLEX $0.560$ $0.370$ 0.343 (+14.7%) $0.482$
@@
Qwen2.5-Coder-7B-SFTNT $0.537$ $0.310$ $0.273$ $0.443$
Qwen2.5-Coder-7B-SFT $0.573$ $0.40$ $0.294$ $0.494$
Qwen2.5-Coder-7B-RLEX $0.619$ $0.463$ 0.406 (+11.2%) $0.552$
Qwen2.5-Coder-7B-RLQATCH 0.628 (+8%) 0.482 (+9.4%) $0.385$ 0.561 (+8.5%)